Calculate the resultant displacement and the direction of a motorbike traveling 40 km to south then 70 km to east and finally 50 km to northwest.


1
Expert's answer
2021-02-09T06:54:00-0500

Let's add up the components of the three vectors and find the components of the resultant displacement:


"R_x=70\ km+50\ km\cdot cos(90^{\circ}+45^{\circ})=34.64\ km,""R_y=-40\ km+50\ km\cdot sin(90^{\circ}+45^{\circ})=-4.64\ km."

Then, we can find the magnitude of the resultant displacement from the Pythagorean theorem:


"R=\sqrt{R_x^2+R_y^2}=\sqrt{(34.64\ km)^2+(-4.64\ km)^2}=34.95\ km."

We can find the direction of the resultant displacement from the geometry:


"\theta=cos^{-1}(\dfrac{R_x}{R})=cos^{-1}(\dfrac{34.64\ km}{34.95\ km})=7.64^{\circ}."


In order to find correct angle we must subtruct the obtained angle from "360^{\circ}":


"\theta=360^{\circ}-7.64^{\circ}=346.1^{\circ}."


The resultant displacement has the magnitude 32.65 km and direction "352.36^{\circ}"(counted conterclockwise from the "x"-axis).